# Choosing observables that capture critical slowing down before tipping points: A Fokker-Planck operator approach

### v1

@@ IV Discussion
Here we propose a method to obtain observables for detecting CSD before TPs in multi-stable systems driven by low noise via a data-driven approximation of the FP operator. While its first $k$ eigenfunctions are very slowly decaying modes related to rare noise-induced escape between metastable states [46], the subsequent eigenfunctions describe probability density patterns in phase space that relax slowest towards the quasi-stationary distribution centered around the metastable states. In the context of TP, where a base state loses stability and the system transitions to an alternative state, we can consider without loss of generality the bistable situation with $k=1$. Further, we consider bifurcation-induced tipping [28], where the system prior to the TP is only observed in the basin of the base state and the contribution of the eigenfunction $\psi_{1}$ remains quasi-constant (until infinitesimally close to the TP). In this case, the first non-trivial eigenfunction that can be observed in data is $\psi_{2}$, which describes the slowest relaxation mode towards the quasi-stationary distribution around the base state within its basin of attraction. As the TP is approached, due to CSD this mode will eventually represent the slowing relaxation along the critical d.o.f.
@@
We suggest to approximate the corresponding eigenfunction of the backward FP operator by the first diffusion coordinate $\xi_{1}$, obtained as scaled eigenvector of the DM Markov matrix. From $\xi_{1}$, one can obtain a physical observable - e.g. by projecting onto patterns obtained as average over data points that extremize $\xi_{1}$ that shows a monotonic increase from the base state along the critical d.o.f towards the edge state. With several examples of low-dimensional bi-stable models we demonstrated that such an observable shows increases in noise-driven fluctuations that provide robust statistical EWS of the CSD associated with the impending TP. We also showed that measuring CSD in the correct observable is crucial when attempting to predict the time of tipping by extrapolating the scaling of variance or autocorrelation of a scalar time series based on the SNB normal form. We furthermore showed that the method can be applied successfully to simulation data from a high-dimensional global ocean model that features a TP of the AMOC. A critical mode was extracted that is in excellent agreement with the mode that would be expected from knowledge of the edge state [45, 15], and a scalar observable was derived from $\xi_{1}$ that shows highly significant increases in variance that are useful as statistical EWS.
@@
A general caveat for this and other methods aiming to measure CSD from high-dimensional systems is that the critical mode may only emerge as the first subdominant eigenmode when already very close to the TP. If there are competing non-critical modes that are very slow, it may not be possible to identify the critical mode, since one would need a very long observational time horizon while being close enough to the TP where the critical mode is finally the slowest. Indeed, it was seen for the Veros model that the correct physical mode takes on the role of $\xi_{1}$, but only as the system was arguably quite close to the TP. Unless one knows beforehand which d.o.f should be measured, for instance by knowledge of the edge state [15] or robust physical considerations, it may only be possible to issue a reliable warning when very close to the TP, and potentially only after the probability of noise-induced transitions has become substantial.
@@
There are other previously proposed dimensionality reduction methods aiming to extract a scalar observable that can be used to detect CSD. These include variance-based techniques [8, 47, 48, 49], where the first principle component is identified by empirical orthogonal functions (EOF) - as originally proposed to obtain the critical mode for EWS [8] - or principal oscillation patterns (POP). An autocorrelation-based method has also been proposed, where the directions of maximum variance of the first differences of multivariate time series are found, which gives components of high autocorrelation that should indicate directions of lowest resilience [50]. Other methods look for a SNB in the full set of (observed) variables via the eigenvalues of a reconstructed Jacobian, which is determined by fitting a multivariate autoregressive model [51] or a multidimensional Langevin equation [52].
@@
Our approach is distinguished by combining several attributes. It yields an observable derived from the first subdominant backward FP eigenfunction that is designed to represent the critical mode displaying CSD, based on the flattening of the quasipotential expected for a broader class of TPs. This mirrors the reasoning for a natural tipping observable recently proposed in [29], and it is also supported by other operator-theoretic work on the topic [53, 54, 55, 56, 57]. The specific DM algorithm that we propose to use is a non-linear dimensionality reduction method and thus allows for observables to be non-linear functions of the state variables. It can be deployed for relatively high dimensional systems, since the quality of its approximation for a given sample size is not dependent on the full state space dimension, but on the intrinsic dimension of the underlying data manifold, which may be much lower. An equal time spacing of data points, or any time ordering at all, is not required. The method furthermore allows for a qualitative assessment of changes in the dominant physical modes, by observing the functional relation between eigenfunctions, and the relation of eigenfunctions and physical variables. This is useful for determining an emerging time scale separation before the TP, and it may help to prevent false positives. It would be interesting in future work to compare our approach to the abovementioned methods.
@@
Future work should address two shortcomings. First, a modification of the method that is viable in higher dimensions and able to reconstruct the correct backward FP eigendecomposition also including non-gradient terms is desirable. This may broaden its applicability and further improve the significance of obtained observables for EWS. Second, in many cases of real-world relevance, such as tipping of the polar ice sheets, the change in the control parameter is fast compared to the critical relaxation mode $\psi_{2}$ and perhaps many other modes. This means that the system is not in a quasi-stationary state, as was assumed here, and it is likely that the critical mode is not displayed before crossing the TP. An extended framework based on non-autonomous dynamical systems theory can hopefully yield useful insights on whether EWS before the de-facto TP exist in this case.
@@ Acknowledgements.
We thank R. Nuterman and the Danish Center for Climate Computing for supporting the simulations with the Veros ocean model. J.L. has received support from Danmarks Frie Forskningsfond (grant no. 2032-00346B) and research grant no. VIL59164 from VILLUM FONDEN.
Appendix A Equations for the five-box ocean model
@@
In this appendix, the equations and parameter values of the five-box ocean model, originally published in [41], are described. The boxes, labelled by $X={N,S,T,IP,B}$, are coupled unidirectionally by the thermohaline overturning circulation $q$, and bi-directionally by the wind-driven circulation. The dynamical equation for box $B$ can be eliminated by salt conservation. The remaining boxes forced by an atmospheric freshwater flux $F_{X}$ multiplied by the reference salinity $S_{0}=0.035$, which is then modulated by $HA_{X}$ to emulate the effect of climate change, where $H$ is the control parameter. The varying strength of the overturning $q$ is proportional to the density difference in the northern and southern boxes, and the temperatures $T_{X}$ are fixed everywhere except in the northern box, where it is assumed that $T_{N}=\mu q+T_{0}$, with a global reference temperature $T_{0}$. This yields
@@
Table A1: Parameter values used for the five-box model, adapted from the FAMOUSA1xCO2 calibration in [41]. $\alpha=0.12$ (thermal coefficient) and $\beta=790$ (haline coefficient) define a linear equation of state for the density of sea water. $V_{i}$ is box volume, $F_{i}$ the freshwater fluxes, $T$ are temperatures, $K_{i}$ are wind fluxes and $A_{i}$ determine the distribution of freshwater forcing. $\eta$ is a mixing parameter between the S and B boxes, $\gamma$ determines the proportion of water which takes the cold-water path, $\lambda$ and $\mu$ are constants. Subscripts indicate box labels, $i\in\{N,T,S,IP,B\}$, and ’00’ indicates a global reference value.
@@
$V_{\rm{N}}(m^{3})$ 3.683 ×$\times 10^{16}$ $F_{\rm{N}}(m^{3}s^{-1})$ 0.375 ×$\times 10^{6}$
$V_{\rm{T}}(m^{3})$ 5.151 ×$\times 10^{16}$ $F_{\rm{T}}(m^{3}s^{-1})$0.723 ×$\times 10^{6}$
$V_{\rm{S}}(m^{3})$ 10.28 ×$\times 10^{16}$ $F_{\rm{S}}(m^{3}s^{-1})$ 1.014 ×$\times 10^{6}$
$V_{\rm{IP}}(m^{3})$ 21.29 ×$\times 10^{16}$ $F_{\rm{IP}}(m^{3}s^{-1})$0.666 ×$\times 10^{6}$
$V_{\rm{B}}(m^{3})$ 88.12 ×$\times 10^{16}$ $F_{\rm{B}}(m^{3}s^{-1})$ 0
$A_{\rm{N}}$ 0.194 $\eta(m^{3}s^{-1})$ 66.061 ×$\times 10^{6}$
@@
$A_{\rm{S}}$0.226 $\lambda(m^{6}kg^{-1}s^{-1})$ 2.66 ×$\times 10^{7}$
$A_{\rm{IP}}$0.565 $\mu(^{o}Cm^{-3}s)$ 7.0 ×$\times 10^{-8}$
$K_{\rm{N}}(m^{3}s^{-1})$ 5.439 ×$\times 10^{6}$ $T_{\rm{S}}(^{o}C)$ 5.571
$K_{\rm{S}}(m^{3}s^{-1})$ 3.760 ×$\times 10^{6}$ $T_{\rm{0}}(^{o}C)$ 3.26
$K_{\rm{IP}}(m^{3}s^{-1})$ 89.778 ×$\times 10^{6}$

### v2

@@ IV Discussion
Here we propose a method to obtain observables for detecting CSD before bifurcation-induced TPs [28] in multi-stable systems via a data-driven approximation of the FP operator. The scenario of bifurcation-induced tipping implies low noise, where the first $k$ FP eigenfunctions $\psi_{k}$ are very slowly decaying modes related to rare noise-induced escape between metastable states [54]. The subsequent eigenfunctions describe probability density patterns in phase space that relax slowest towards a quasi-stationary distribution centered around a metastable state. We consider without loss of generality the bistable situation with $k=1$, where at the TP a base state loses stability and the system transitions to its alternative state. In practical applications, the system prior to the TP is only observed in the basin of the base state and the contribution of $\psi_{1}$ remains quasi-constant (until infinitesimally close to the TP). In this case, the first non-trivial eigenfunction that can be observed in data is $\psi_{2}$, which describes the slowest relaxation mode towards the quasi-stationary distribution around the base state within its basin of attraction. As the TP is approached, due to CSD this mode will eventually represent the slowing relaxation along the critical d.o.f.
@@
We suggest to approximate the corresponding eigenfunction of the backward FP operator by the first diffusion coordinate $\xi_{1}$, obtained as scaled eigenvector of the DM Markov matrix. From $\xi_{1}$, one can obtain a physical observable - e.g. by projecting onto patterns obtained as average over data points that extremize $\xi_{1}$ - that shows a monotonic increase from the base state along the critical d.o.f towards the edge state. With several examples of low-dimensional bi-stable models we demonstrated that such an observable shows increases in noise-driven fluctuations that provide robust statistical EWS of the CSD associated with the impending TP. We also showed that measuring CSD in the correct observable is crucial when attempting to predict the time of tipping by extrapolating the scaling of variance or autocorrelation of a scalar time series based on the SNB normal form. The method was successfully applied to simulation data from a high-dimensional global ocean model that features a TP of the AMOC. A critical mode was extracted that is in excellent agreement with the mode that would be expected from knowledge of the edge state [53, 15], and a scalar observable was derived from $\xi_{1}$ that shows highly significant increases in variance that are useful as statistical EWS.
@@
A general caveat for this and other methods aiming to measure CSD from high-dimensional systems is that the critical mode may only emerge as the first subdominant eigenmode when already very close to the TP. If there are competing non-critical modes that are very slow, it may not be possible to identify the critical mode, since one would need a very long observational time horizon while being close enough to the TP where the critical mode is finally the slowest. Indeed, it was seen for the Veros model that the correct physical mode takes on the role of $\xi_{1}$, but only as the system was quite close to the TP. Unless one knows beforehand which d.o.f should be measured, for instance by knowledge of the edge state [15] or robust physical considerations, it may only be possible to issue a reliable warning when very close to the TP, and potentially only after the probability of noise-induced transitions has become substantial.
@@
There are other previously proposed dimensionality reduction methods aiming to extract a scalar observable that can be used to detect CSD. These include variance-based techniques [8, 55, 56, 57], where the first principle component is identified by empirical orthogonal functions - as originally proposed to obtain the critical mode for EWS [8] - or principal oscillation patterns. An autocorrelation-based method has also been proposed, where the directions of maximum variance of the first differences of multivariate time series are found, which gives components of high autocorrelation that should indicate directions of lowest resilience [58]. Other methods look for a SNB in the full set of (observed) variables via the eigenvalues of a reconstructed Jacobian, which is determined by fitting a multivariate autoregressive model [59] or a multidimensional Langevin equation [44].
@@
Our approach is distinguished by combining several attributes. It yields an observable derived from the first subdominant backward FP eigenfunction that is designed to represent the critical mode displaying CSD, based on the flattening of the quasipotential expected for a broader class of TPs. This mirrors the reasoning for a natural tipping observable recently proposed in [29], and it is also supported by other operator-theoretic work on the topic [60, 61, 62, 63, 64]. The DM algorithm that we propose to use is a non-linear dimensionality reduction method and thus allows for observables to be non-linear functions of the state variables. It can be deployed for relatively high dimensional systems, since the quality of its approximation for a given sample size is not dependent on the full state space dimension, but on the intrinsic dimension of the underlying data manifold, which may be much lower. An equal time spacing of data points, or any time ordering at all, is not required, thus increasing its applicability for instance to paleoclimate data [65]. The method furthermore allows for a qualitative assessment of changes in the dominant physical modes, by observing the functional relation between eigenfunctions, and the relation of eigenfunctions and physical variables. This is useful for determining an emerging time scale separation before the TP, and it may help to prevent false positives for instance due to changing amplitude or correlation of the driving noise. Estimating the eigenfunctions with the generalized DM method presented in [36] also allows treating multiplicative noise and non-gradient dynamics, which can alleviate the issues that unknown changes in oscillatory modes cause for statistical EWS [52]. This method gives equivalent results when applied to the ocean model (see App. A).
@@
Future work should address how systems driven far away from a steady state need to be treated. In many cases of real-world relevance, such as tipping of the polar ice sheets, the change in the control parameter is fast compared to the critical relaxation mode $\psi_{2}$ and perhaps many other modes. This means that the system is not in a quasi-stationary state, as was assumed here, and it is likely that the critical mode is not displayed before crossing the TP. An extended framework based on non-autonomous dynamical systems theory can hopefully yield useful insights on whether EWS before the de-facto TP exist in this case.
@@ Acknowledgements.
GAG acknowledges valuable discussions with Matheus Mankato de Castro and Caroline Wormell. We thank R. Nuterman and the Danish Center for Climate Computing for supporting the simulations with the Veros ocean model. J.L. has received support from Danmarks Frie Forskningsfond (grant no. 2032-00346B) and research grant no. VIL59164 from VILLUM FONDEN.
Appendix A Eigenfunctions in non-gradient systems with non-additive noise estimated by local Kernel diffusion maps
Here we present how the Fokker-Planck and backward Kolmogorov eigenfunctions can be estimated without assuming gradient dynamics or additive noise, and demonstrate that similar results are obtained for data from the Veros model as with the standard diffusion map method presented in the main text. It can be achieved by the method introduced in [36], which generalizes diffusion maps to apply to SDEs of the form (1). In addition to the isotropic Kernel (11), an anisotropic Kernel with a second bandwidth $\epsilon_{l}$, as introduced in [66], is used:
$K_{\epsilon_{l}}(\mathbf{x},\mathbf{y})=\exp\left(-(4\epsilon_{l})^{-1}\left(\mathbf{x}-\mathbf{y}+\epsilon_{l}\hat{b}(\mathbf{x})\right)^{\mathsf{T}}\left(\hat{A}(\mathbf{x})\right)^{-1}\left(\mathbf{x}-\mathbf{y}+\epsilon_{l}\hat{b}(\mathbf{x})\right)\right).$ (23)
This requires estimates of the drift $\hat{b}(\mathbf{x})$ and diffusion matrix $\hat{A}(\mathbf{x})$ locally at each data point $\mathbf{x}$. To do so we use Kramers-Moyal expansion estimates
$\hat{b}(\mathbf{x})$ $=\lim_{\tau\to 0}\frac{1}{\tau}\mathbf{E}[X_{\tau}-X_{0}|X_{0}=\mathbf{x}]$ (24)
$\hat{A}(\mathbf{x})$ $=\lim_{\tau\to 0}\frac{1}{2\tau}\mathbf{Cov}[X_{\tau}-X_{0}|X_{0}=\mathbf{x}].$ (25)
Ideally, in order to calculate expectation values, one has many realizations with short integration time $\tau$ at each data point as initial condition $\mathbf{x}$. But in the case of the Veros model, we only have one long realization (for each parameter value), and thus we estimate drift and diffusion at each data point $\mathbf{x}$ by computing expectation and covariance over the set of increments $\{Y_{n_{i}+\tau}-Y_{n_{i}}\}_{i=1}^{n}$, where $\{Y_{n_{i}}\}_{i=1}^{n}$ are the $n$ nearest neighbors of $\mathbf{x}$. While we cannot take the limit $\tau\to 0$, the sample time $\tau$ of 5 years is short compared to the long time scales of the deep ocean.
After computing the anisotropic Kernel $(K_{\epsilon_{l}})_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$ for all pairs of data points, it is normalized using the row sum of the isotropic Kernel $p_{\epsilon}(\mathbf{x}_{i})=\sum_{j=1}^{N}K(\mathbf{x}_{i},\mathbf{x}_{j})$, i.e., a Kernel density estimate of the data sample. Thus, with the diagonal matrix $(D_{\epsilon})_{ii}=p_{\epsilon}^{-1}(\mathbf{x}_{i})$, form $\tilde{K}_{\epsilon_{l}}=K_{\epsilon_{l}}D_{\epsilon}$. This normalization removes additional drift terms in the reconstructed operator, which are induced by the non-uniform sampling density. Finally, to conserve probability, the matrix is normalized by its row sums, i.e., $L_{\epsilon_{l}}=D_{\epsilon,\epsilon_{l}}^{-1}\tilde{K}_{\epsilon_{l}}$ with the diagonal matrix $(D_{\epsilon,\epsilon_{l}})_{ii}=\sum_{j=1}^{N}(\tilde{K}_{\epsilon_{l}})_{ij}$. As shown in [36], in the limits $N\to\infty$, $\epsilon\to 0$, and $\epsilon_{l}\to 0$ the matrix $(L_{\epsilon_{l}}-I)/\epsilon_{l}$ converges at every point to the generator (backward FP operator) $\mathcal{L}^{*}$ of the general SDE (1), and we can approximate $\{\phi_{n}\}$ by the eigenfunctions of $L_{\epsilon_{l}}$. The forward FP operator can be constructed in a similar fashion [36].
Figure 14: a-d Anomaly maps of deep ocean (vertical average below 1000m) temperature (a,b) and salinity (c,d) describing the first subdominant spatiotemporal mode $\xi_{1}$ of the Veros model obtained by the local Kernel DM algorithm [36] at the control parameter values $F=0.3472$ (a,c) and $F=0.3557$ (b,d). The modes are constructed in the same way as described in Sec. III.5. e,f Spectrum of the local Kernel diffusion map (higher eigenvalues discarded). Note that the reported eigenvalues of the matrix $L_{\epsilon_{l}}$ correspond to $e^{\lambda_{n}}$ of the approximated generator.
For the analysis of the Veros data, we chose $\epsilon_{l}/\tau=5$ and $n=100$. As explained in [36], the bandwidth may be interpreted as a characteristic time scale, and is chosen here to match the average displacement in phase space after one sample time $\tau$. The estimate of the inverse diffusion matrix is regularized by replacing $\hat{A}(\mathbf{x})$ with $\hat{A}(\mathbf{x})+\eta I$ in (23). A large regularization parameter of $\eta=0.3$ had to be chosen due to the small sample size. In Fig. 14 the results are summarized. We focus on the parameter very close to the bifurcation ($F=0.3557$), as well as one further away ($F=0.3472$), just as presented in Fig. 13. The evolution of the eigenfunctions is very similar to the isotropic diffusion map method. Since the generalized method also picks up non-gradient dynamics, there are oscillatory modes, but they relax more quickly and thus appear higher up in the spectrum (Fig. 14e,f). The initially dominant mode of cold excursions initiated in the Southern ocean is picked up more clearly with the generalized method (Fig. 14a,c). The critical mode close to the TP (Fig. 14b,d) is again very similar to the fingerprint of the edge state (Fig. 13c,f) and thus allows strong EWS by projecting time series data onto it.
Appendix B Equations for the five-box ocean model
@@
In this appendix, the equation and parameter values of the five-box ocean model, originally published in [49], are described. The boxes, labelled by X=N,S,T,IP,BX={N,S,T,IP,B}, are coupled unidirectionally by the thermohaline overturning circulation $q$, and bi-directionally by the wind-driven circulation. The dynamical equation for box $B$ can be eliminated by salt conservation. The remaining boxes forced by an atmospheric freshwater flux $F_{X}$ multiplied by the reference salinity $S_{0}=0.035$, which is then modulated by $HA_{X}$ to emulate the effect of climate change, where $H$ is the control parameter. The varying strength of the overturning $q$ is proportional to the density difference in the northern and southern boxes, and the temperatures $T_{X}$ are fixed everywhere except in the northern box, where it is assumed that $T_{N}=\mu q+T_{0}$, with a global reference temperature $T_{0}$. This yields
@@
Table A1: Parameter values used for the five-box model, adapted from the FAMOUSA1xCO2 calibration in [49]. $\alpha=0.12$ (thermal coefficient) and $\beta=790$ (haline coefficient) define a linear equation of state for the density of sea water. $V_{i}$ is box volume, $F_{i}$ the freshwater fluxes, $T$ are temperatures, $K_{i}$ are wind fluxes and $A_{i}$ determine the distribution of freshwater forcing. $\eta$ is a mixing parameter between the S and B boxes, $\gamma$ determines the proportion of water which takes the cold-water path, $\lambda$ and $\mu$ are constants. Subscripts indicate box labels, $i\in\{N,T,S,IP,B\}$, and ’0’ indicates a global reference value.
@@
$V_{\rm{N}}(m^{3})$ 3.683 $\times 10^{16}$ $F_{\rm{N}}(m^{3}s^{-1})$ 0.375 $\times 10^{6}$
$V_{\rm{T}}(m^{3})$ 5.151 $\times 10^{16}$ $F_{\rm{T}}(m^{3}s^{-1})$ -0.723 $\times 10^{6}$
$V_{\rm{S}}(m^{3})$ 10.28 $\times 10^{16}$ $F_{\rm{S}}(m^{3}s^{-1})$ 1.014 $\times 10^{6}$
$V_{\rm{IP}}(m^{3})$ 21.29 $\times 10^{16}$ $F_{\rm{IP}}(m^{3}s^{-1})$ -0.666 $\times 10^{6}$
$V_{\rm{B}}(m^{3})$ 88.12 $\times 10^{16}$ $F_{\rm{B}}(m^{3}s^{-1})$ 0
$A_{\rm{N}}$ 0.194 $\eta(m^{3}s^{-1})$ 66.061 $\times 10^{6}$
@@
$A_{\rm{S}}$ -0.226 $\lambda(m^{6}kg^{-1}s^{-1})$ 2.66 $\times 10^{7}$
$A_{\rm{IP}}$ -0.565 $\mu(^{o}Cm^{-3}s)$ 7.0 $\times 10^{-8}$
$K_{\rm{N}}(m^{3}s^{-1})$ 5.439 $\times 10^{6}$ $T_{\rm{S}}(^{o}C)$ 5.571
$K_{\rm{S}}(m^{3}s^{-1})$ 3.760 $\times 10^{6}$ $T_{\rm{0}}(^{o}C)$ 3.26
$K_{\rm{IP}}(m^{3}s^{-1})$ 89.778 $\times 10^{6}$